SpaceRef

SpaceRef


ISS On-Orbit Status 19 June 2002

Status Report From: NASA HQ
Posted: Wednesday, June 19, 2002

All ISS systems continue to function nominally, except as noted previously or below.  The new station residents have settled in and are doing fine.

Another "light-duty" day aboard the ISS:  5.5 hours of work for each crewmember plus one hour of  ISS familiarization/orientation.

CDR Valery Korzun and FE-2 Sergej Treschev took turns as subject and CMO/assistant in completing a session with the Russian PZE-MO-5 MedOps protocol of cardiovascular evaluation during graded exercises on the cycle ergometer over a Russian ground site (12-15 min.) each. [Using the Gamma-1M ECG equipment with biomed harness, skin electrodes, and a blood pressure and rheoplethysmograph cuff wired to the cycle ergometer's instrumentation panels, each cosmonaut worked the pedals after a prescribed program at load settings of 125, 150, and 175 watts for three minutes each.  All measurements were recorded and telemetered to MCC-M, from where the workout was controlled.]

FE-1 Peggy Whitson had her first science day today, working on the BSTC (biotechnology specimen temperature controller) payload.  To initiate part 1 of the experiment, she removed the four QTCMA bioreactors (#3, 4, 5, 6) from the ARCTIC1 freezer and placed them in the BSTC. In part 2, she then inoculated selected module assemblies with cells for culturing in nutrients, and fixated (preserved) others.  [BSTC is part of NASA/JSC's Cellular Biotechnology Program which develops ground-based and space bioreactor technology to support investigations in cell biology and tissue engineering by providing the proper thermal and gas environment plus employing the weightless environment to form three-dimensional, functional tissue equivalents. The BSTC houses stationary bioreactors called QTCMAs (quad tissue culture module assemblies) which it maintains at a specified temperature in a controlled atmosphere. Growth, morphology and function of mammalian cells in the BSTC are monitored by Whitson by means of pH and chemistry analysis of media samples and postflight analysis of "fixed" cell samples. A gas supply module (GSM) provides carbon dioxide gas for purging the incubation chambers after each daily session, and the BCSS (biotechnology cell science stowage), which includes a supercold cryo-dewar, supplies the consumables and tools.  Peggy's task during the duration of the experiment is to analyze the media in each QTCMA, replace them with fresh media (using syringes, gloves and a face mask), preserve cultures, take images with the Kodak digital still camera, and purge with CO2 at the end. Each session takes about three hours minimum].

After yesterday's freezer/refrigerator swap of the ARCTIC (advanced thermoelectric refrigerator/freezer) system, the water flow rate in the new configuration is still lower than expected, preventing ARCTIC1 to reach the desired temperature of -20 degC.  FE-1 Whitson was asked to check out four QDs (quick disconnects) of the water hoses and also look for anything abnormal in the lines, such as kinks, with particular attention to ARCTIC2.

Korzun and Treschev continued packing and stowing discarded equipment, trash and waste containers in Progress-257 (7P).  Part of the demanding task is the photo and video recording of activities for evaluation by the ground prior to Progress undock.  The photo/video recordings of today were downlinked via U.S. assets (Ku-band). [For proper stability and control of the fully automated spacecraft, operating on pre-stored command algorithms, it is crucial that center of gravity and mass moments of inertia of the ship remain within prescribed limits.  Image records are also helpful for maintaining the IMS (inventory management system).]

Peggy Whitson's live interactive TV downlink to the Senate Subcommittee Hearing on NASA's relationship to education was cancelled today since the TDRS satellite for this time slot was required for the late Shuttle landing at Edwards Air Force Base.  However, a video message had been downlinked and taped on the ground on 6/17 for the event.

CDR Korzun performed some additional component replacements in the SM's ASU toilet compartment, changing out the pre-treat container and associated hose.

FE-1 Whitson completed installation of new HRM (heart rate monitor) software in the MEC (medical equipment computer) and then set up the s/w application with specific crew data (name, date of birth, height in cm).  [Since she was unable yesterday to locate the bootable floppy for the MEC laptop, required for the s/w installation, she first formatted a new systems disk using a CD-ROM in the laptop, for which MCC-H had uplinked instructions.]

Peggy also completed her first U.S. MedOps crew nutrition assessment by filling out the standard food frequency questionnaire (FFQ) on the MEC desktop.

MCC-Moscow (TsUP) ran remote tests of the ASN satellite navigation system, GTS global timing system and program logic control device (UPLU) with the Russian "Wiener" power laptop OpsLAN connection (BSMM), a part of the onboard equipment control system (SUBA).  No crew action was required.  [SUBA is used to control, monitor, and diagnose SM onboard systems status by employing sensor output signals, command radio link SM functional outputs and other units.]

In preparation for the Progress-257 undocking on 6/25, TsUP plans a Service Module (SM) fuel line purge for 6/21 (Friday).  The fuel will be discharged from a vent located on the cylindrical portion of the unpressurized SM Assembly Compartment (AO), and peripheral portions of the gas-liquid phase discharge may be observed from a window in the crew cabin (resembling a stream of snow flakes, with uneven intensities). The crew was advised to capture it on video with the LIV camcorder, starting at 3:25 am EDT.  Total duration of the purge is 13 minutes, and the visible portion lasts about 5 min.

STS-111/Endeavour landed (finally) at Edwards AFB at 1:58 pm EDT, after again having being waved off at KSC for both of today's east coast opportunities.  Total mission duration was 13d 20h 35m (217 orbits).  Welcome back, Yuri, Carl and Dan!   [This brings Carl's and Dan's U.S. record in space to 196d 20h 39m (189 days on board ISS).]

Today's target areas for the Crew Earth Observation (CEO) program were E. Mediterranean Dust and Smog (this pass tracked through the target area from northeastern Egypt to southeastern Turkey. With high pressure and light winds, crew was to try for oblique views to the right of track of aerosol accumulation over the urbanized Nile Delta and eastern Mediterranean coast), European Smog (as the ISS tracked northeastward from Tunisia to the Adriatic Sea, the crew was to try for oblique views of smog over the coastal waters of both eastern and western Italy, to the left of track), High Central Andean Glaciers (Sun was low this pass, but cloud cover was expected to be much lower than usual.  Crew was to use the long lenses on the ESC to try for near-nadir views of the details of the small, isolated ice fields and glaciers on the crest of the Andes), Eastern United States (this pass tracked northeastward from just west of the Appalachians to northern New England.  High pressure and light winds to the right of track should afford oblique views of smog over the US eastern seaboard), Gulf of St. Lawrence (while ice season ended for the Gulf of St Lawrence early last month, opportunities to sight icebergs are still possible over the Labrador Sea, NE of Newfoundland. As ISS crossed the east coast, the crew was to look northward [left of track] for icebergs moving southeastward into shipping lanes on the Labrador Current.  On the second and northernmost pass of the day over this target, the interest was on oblique views to the left of track of the remaining sea ice swirling along the rugged coast of Labrador.  On the third and final pass over this target area, light was getting low.  However, the crew may still have spotted icebergs in the open waters northeast of Newfoundland).

CEO images can be viewed at the website http://eol.jsc.nasa.gov

U.S. and Russian Segment Status (as of 1:12 am EDT):

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):
Elektron O2 generator is powered On (32-amp mode). Vozdukh CO2 scrubber is ON in MANUAL cycle mode #5 (vacuum pump failed).  U.S. CDRA CO2 scrubber is Off (failed).  BMP Harmful Impurities unit: Absorbent bed #1 in Purify mode, bed #2 in Regeneration mode.

SM Working Compartment:  Pressure (mmHg) -- 764, temperature (deg C) -- 26.3, ppO2 (mmHg) -- 155.3, ppCO2 (mmHg) -- 2.6.
SM Transfer Compartment:  Pressure (mmHg) -- 762, temperature (deg C) -- 19.9.
FGB Cabin:  Pressure (mmHg) -- 756, temperature (deg C) -- 20.7.
Node:  Pressure (mmHg) -- 755.34, temperature (deg C) -- 22.1 (shell); ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
U.S. Lab:  Pressure (mmHg) -- 757.64, temperature (deg C) -- 25.1, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a;
Joint Airlock (Equip. Lock):  Pressure (mmHg) -- 757.54, temperature (deg C) -- 22.6; shell heater temp (deg C) -- 23.1, ppO2 (mmHg) -- n/a; ppCO2 (mmHg) -- n/a.
PMA-1:  Shell heater temp (deg C) -- 22.4
PMA-2:  Shell heater temp (deg C) -- 17.4

(Note: Partial pressures ppO2 and ppCO2 in U.S. segment [USOS] not available because MCA [major constituent analyzer] is failed and in Extended Life mode [= a state that preserves mass spectrometer vacuum but produces no pp data]). MSA (mass spectrometer assembly) and VGA (verification gas assembly) have been removed for return to Earth.

Electrical Power Systems (EPS):
Beta Gimbal Assembly (BGA) 2B in Autotrack; BGA 4B in Autotrack mode.
SM batteries: Battery #6 is cycling; all other batteries (7) are in "Partial charge" mode.
FGB: Battery #2 is offline (ROM mode). All other batteries (5) are in "Partial charge" mode.
Plasma Contactor Unit PCU-1 in Standby mode; PCU-2 in Standby mode.

Thermal Control Systems:
Air conditioner SKV-1 is Off. SKV-2 is On.

Command & Data Handling Systems:
C&C-3 MDM is prime, C&C-2 is back-up, and C&C-1 is in standby.
GNC-1 MDM is prime; GNC-2 is Backup.
LA-1, LA-2 and LA-3 MDMs are all operating.
PL-1 MDM is operational; PL-2 MDM is Off.
APS-1 (automated payload switch #1) and APS-2 are both On.
SM Terminal Computer (TVM): 2 redundant lanes (of 3) operational.
SM Central Computer (TsVM): 2 redundant lanes (of 3) operational.

Attitude Source:
State vector -- US GPS (SIGI string 1)
Attitude -- Russian
Angular rates -- US RGA1 (rate gyro assembly 1)

Communications Systems:
All Russian communications & tracking systems are nominal.
S-band is operating nominally.
Ku-band is operating nominally.
Audio subsystem operating nominally.
Video subsystem operating nominally.
MCOR (medium-rate communications outage recorder) is powered Off (due to consistently reaching upper operational temperature during ops).

Robotics:
SSRMS/Canadarm2 at MBS PDGF (mobile base system/power & data grapple fixture), with Keep Alive power on both strings.
RWS (robotics workstations): Cupola is Active, Lab RWS is Off.
SSRMS Prime string Wrist Roll (WR) joint was replaced on UF-2.

ISS Orbit (as of this morning, 6:28am EDT [= epoch]):
Mean altitude -- 393.7 km
Apogee -- 398.4 km
Perigee -- 389.1 km
Period -- 92.4 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0006897
Orbits per 24-hr. day -- 15.58
Altitude decrease -- 200 m (mean) in last 24 hours
Solar Beta Angle --  -16 deg (magnitude increasing)
Revolutions since FGB/Zarya launch (Nov. '98) -- 20438
Current Flight Attitude -- LVLH (local vertical/local horizontal = "earth-fixed": z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7 deg., roll: 0 deg])

For more on ISS orbit and naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.